### v1

@@ 3.1 Alpha filtration
[…]
@@ 3.2 Persistent homology and representative cycle
A useful tool for visualizing persistent homology is the persistence diagram (PD) [25], as shown in Fig. 4. The set of points that record the birth time and death time of $n$-cycles is called the $n$-PD. Each point in a $n$PD is represented as $(b_{i},d_{i})$, corresponding to an $n$-cycle and capturing its birth time $b_{i}$ and death time $d_{i}$. The value $\left|b_{i}-d_{i}\right|$ is defined as the persistence of this cycle. Additionally, consider a filtration that introduces one simplex at a time (an alpha filtration satisfies this condition in practice). We define a simplex $\sigma$ as positive if its addition creates a cycle, thereby giving birth to a new homology class. Otherwise, we define it as negative. For a 2-cycle in an alpha filtration, the positive simplex is a 2-simplex, while the negative simplex is a 3-simplex. Actually, for a point in $n$-PD and its corresponding $n$-cycle, the birth time indicates when the positive simplex appears in the filtration, while the death time indicates when the negative simplex appears in the filtration.
@@
[…]
@@ 4.1 Topological understanding of the point cloud
To identify significant points in 2-PD, we initially project the points in 2-PD onto the line $y=-x$. The distance from these projected points to the origin (0,0)(0,0)( 0 , 0 ) serves as a proxy for the persistence of the corresponding points in the 2-PD. The projected points are then divided into two classes using the $k$-means clustering algorithm (in our method $k=2$, since in the analysis of the persistence diagram, the common practice is to divide the points inside into two classes, i.e. significant points and noise points) [34]. Points belonging to the class with greater persistence are regarded as significant, and each significant point corresponds to a closed surface that is to be reconstructed from the point cloud. Fig. 5 illustrates this clustering process of a PD.
@@ 4.2 Computing representative 2-cycles
[…]
@@
[…]
$\partial[ABCD]=[BCD]+[ACD]+[ABD]+[ABC],$
and for $[BCD]\in K(r_{4})-K(r_{1})$,
$[BCD]^{*}(\partial[ABCD])=[BCD]^{*}([BCD])=1\neq 0.$
[…]
@@ 4.3 Removing non-manifold vertices and edges
Suppose we have obtained a persistent volume $vol$ and the volume-optimal cycle $c$ corresponding to a significant 2-PD point. For each vertex $v$ in $c$, we first check the neighboring faces $F(v)$ of $v$. Starting from one face $F_{1}\in F(v)$, we identify another face $F_{2}\in F(v)$ that shares a common edge with $F_{1}$. This search process is repeated until $F_{1}$ is reencountered. If all faces in $F(v)$ are visited during this process, $v$ is classified as a normal vertex; otherwise, it is considered a non-manifold vertex. After finding all non-manifold vertices, we remove all 3-simplices in $vol$ adjacent to at least one non-manifold vertex, thus deriving an updated $vol$ and $c$.
Next, we eliminate non-manifold edges. Since the volume-optimal cycle $c$ is the boundary of a combination of 3-simplices, we define non-manifold edges in $c$ as those connected to more than two faces. After identifying these edges and removing all their neighboring 3-simplices in $vol$, the cleaned persistent volume $vol^{new}$ will not contain non-manifold vertices or edges. Hence, we finally compute $\hat{c}=\partial(vol^{new})$ again. The resulting $\hat{c}$ is now a representative 2-cycle without non-manifold vertices and edges, which is feasible for further optimization steps.
@@ 4.4 Loop subdivision fitting by LSPIA
[…]

### v2

@@ 3.1 Alpha filtration
where $\|\cdot\|$ denotes the Euclidean norm in $\mathbb{R}^{n}$. The Delaunay complex generated by $P$ is given by
@@ 3.2 Persistent homology and representative cycle
A useful tool for visualizing persistent homology is the persistence diagram (PD) [25], as shown in Fig. 4. The set of points that record the birth time and death time of $n$-cycles is called the $n$-PD. Each point in a $n$ -PD is represented as $(b_{i},d_{i})$, corresponding to an $n$-cycle and capturing its birth time $b_{i}$ and death time $d_{i}$. The value $\left|b_{i}-d_{i}\right|$ is defined as the persistence of this cycle. Additionally, consider a filtration that introduces one simplex at a time (an alpha filtration satisfies this condition in practice). We define a simplex $\sigma$ as positive if its addition creates a cycle, thereby giving birth to a new homology class. Otherwise, we define it as negative. For a 2-cycle in an alpha filtration, the positive simplex is a 2-simplex, while the negative simplex is a 3-simplex. Actually, for a point in $n$-PD and its corresponding $n$-cycle, the birth time indicates when the positive simplex appears in the filtration, while the death time indicates when the negative simplex appears in the filtration.
@@
For each homology group $H_{n}(K)$, suppose its basis is $\left\{[r_{1}],\cdots,[r_{k}]\right\}$, then each $r_{i}\>(i=1,\cdots,k)$ is referred to as a representative $n$-cycle of the homology class $r_{i}]$. In the context of persistent homology, the representative $n$-cycle of a persistent pair $(b,d)$, known as a persistent $n$-cycle [32], is a specific representative $n$-cycle in the filtration $K^{b}\subset K^{b+1}\subset\cdots\subset K^{d}$. For instance, a persistent $n$-cycle associated with $(b,d)$ first emerges at $K^{b}$ and becomes a boundary at $K^{d}$. In our method, we adopt the volume optimal cycle [33] as the representative cycle for each point in the 2-PD.
@@ 4.1 Topological understanding of the point cloud
To identify significant points in 2-PD, we initially project the points in 2-PD onto the line y=xy=-x. The distance from these projected points to the origin $(0,0)$ serves as a proxy for the persistence of the corresponding points in the 2-PD. The projected points are then divided into two classes using the $k$-means clustering algorithm (in our method $k=2$, since in the analysis of the persistence diagram, the common practice is to divide the points inside into two classes, i.e. significant points and noise points) [34]. Points belonging to the class with greater persistence are regarded as significant, and each significant point corresponds to a closed surface that is to be reconstructed from the point cloud. Fig. 5 illustrates this clustering process of a PD.
@@ 4.2 Computing representative 2-cycles
Here we use Fig. 6 as an example of computing volume-optimal cycle. In this filtration, the simplicial complex $K(r_{0})$ contains five 2-simplices: $[ABD]$, $[ACD]$, $[BCE]$, $[BDE]$, $[CDE]$ (shown in red). In $K(r_{1})$, the 2-simplex $[ABC]$ is added, generating a 2-cycle:
@@
Then in $K(r_{2})$, $[BCD]$ is added, while the 3-simplices $[ABCD]$ and $[BCDE]$ (shown in blue) are added in $K(r_{3})$ and $K(r_{4})$, respectively. We observe that in this filtration, there is a persistent pair $(r_{1},r_{4})$, where $[ABC]$ is the positive simplex and $[BCDE]$ is the negative simplex. We now compute the persistent volume and volume-optimal cycle for it. According to the definition, the persistent volume must contain $[BCDE]$, and thus it can be either $[BCDE]$ or $[ABCD]+[BCDE]$. However,
$\partial[BCDE]=[CDE]+[BDE]+[BCE]+[BCD],$
and for $BCD]\in K(r_{4})-K(r_{1})$,
$BCD]^{*}(\partial[BCDE])=[BCD]^{*}([BCD])=1\neq 0.$
Thus $[BCDE]$ does not satisfy the condition (Eq. 2) and is not a persistent volume for $(r_{1},r_{4})$. On the other hand, we can similarly check that $[ABCD]+[BCDE]$ is a persistent volume for $(r_{1},r_{4})$, and its boundary
@@ 4.3 Removing non-manifold vertices and edges
Suppose we have obtained a persistent volume volvol and the volume-optimal cycle $c$ corresponding to a significant 2-PD point. For each vertex $v$ in $c$, we first check the neighboring faces $F(v)$ of $v$. Starting from one face $F_{1}\in F(v)$, we identify another face $F_{2}\in F(v)$ that shares a common edge with $F_{1}$. This search process is repeated until $F_{1}$ is reencountered. If all faces in $F(v)$ are visited during this process, $v$ is classified as a normal vertex; otherwise, it is considered a non-manifold vertex. After finding all non-manifold vertices, we remove all 3-simplices in volvol adjacent to at least one non-manifold vertex, thus deriving an updated volvol and $c$.
Next, we eliminate non-manifold edges. Since the volume-optimal cycle $c$ is the boundary of a combination of 3-simplices, we define non-manifold edges in $c$ as those connected to more than two faces. After identifying these edges and removing all their neighboring 3-simplices in volvol, the cleaned persistent volume $vol^{new}$ will not contain non-manifold vertices or edges. Hence, we finally compute $\hat{c}=\partial(vol^{new})$ again. The resulting $\hat{c}$ is now a representative 2-cycle without non-manifold vertices and edges, which is feasible for further optimization steps.
@@ 4.4 Loop subdivision fitting by LSPIA
Point Cloud $\begin{array}[]{c}\text{ Number of }\\ \text{ Points }\end{array}$ $\begin{array}[]{c}\text{ Number of }\\ \text{ Control Vertices }\end{array}$ $\begin{array}[]{c}\text{ Number of }\\ \text{ Surfaces }\end{array}$ RMS $\begin{array}[]{c}\text{ Time of }\\ \text{ Computing Topology }\end{array}$ $\begin{array}[]{c}\text{ Time of }\\ \text{ Subdivision Fitting }\end{array}$